State the degrees of freedom error in each of the following tests.

(a) A consultant measures job satisfaction in a sample of 14 supervisors, 14 managers, and 14 executives at a local firm.


(b) A researcher tests how nervous public speakers get in front of a small, medium, or large audience. Eight participants are randomly assigned to each group.


(c) A high school counselor has 9 students in each of five classes rate how much they like their teacher.
Answer #1

Ans:

a)

N=14+14+14=42

number of groups,k=3

degrees of freedom (error)=42-3=39

b)

N=3*8=24

number of groups,k=3

degrees of freedom (error)=24-3=21

c)

N=5*9=45

number of groups,k=5

degrees of freedom (error)=45-5=40